Jim is older than Bob. Together, they are 60 years old. If Bob is 8 years younger than Jim, find the age of each.
Answer by FrankM(1040)

J+B = 60
B+8=J
So we substitute:
(B+8)+B=60
2B+8=60
2B=52
B=26
J=34